# Conformal Convolution and Monte Carlo Meta-learners for Predictive Inference of Individual Treatment Effects

### v5

@@ Appendix D The Epsilon Problem: (un)identifiability of ITE predictive distribution
We formally state the identifiability constraint:
@@
Case 3. No identified copula If neither the independence nor a known copula is assumed, then infinitely many different copulas are compatible with the same pair of marginals but induce different distributions of the difference $Y(1)-Y(0)$.
@@ D.1 Prediction intervals and violations of the noise assumption
[…]
@@ E.1 Probabilistic calibration of weighted conformal predictive system
The simplification at ()(*)( ∗ ) can be made because we consider only a special case of weighted exchangeability, the covariate shift, where $w_{1:n}\equiv 1$ and $w_{n+1}((x,y))=w(x)=\frac{d\tilde{P}_{X}(x)}{dP(x)}$. Nevertheless, our findings can be readily applied to the general concept of weighted exchangeability.
@@ E.3 Existence universal consistent CCT- and CMC T-learner
[…]
@@
the outcome variables are independent conditionally on the covariate set $X$. Then, by the definition of convolution and the fact that $ITE=Y(1)-Y(0)$, the convolution of the true conditional cumulative distribution of outcome under treatment $F_{Y(1)|X}$ and outcome under negative control $F_{-Y(0)|X}$ equals the true conditional cumulative ITE distribution,
[…]
@@ G.2 Synthetic experiments
This data generation process entails both Setup A and Setup B. For Setup A, where $\gamma=1$, represents an experiment where the treatment has no effect, while for Setup B, where $\gamma=0$, represents a heterogeneous treatment effect. X is sampled from a uniform distribution, $d$ is the number of covariates, $\tau$ represents the treatment effect, $\gamma$ is the treatment heterogeneity, $\pi$ is the propensity, $\epsilon(0)$ and $\epsilon(1)$ are Gaussian noise terms, and $W$ is the treatment assignment.
@@
$\epsilon(1)\sim\mathcal{N}(0,1)$ (30)
$\epsilon(0)\sim\mathcal{N}(0,1)$ (31)
$Y(0)=\tau_{0}+\epsilon(0)$ (32)
$Y(1)=\tau+\epsilon(1)$ (33)
@@ G.2.2 Nie and Wager synthetic data experiments
Besides the aforementioned experiments with the data-generating processes proposed by Nie and Wager [2021], we also perform a noise dependency impact experiment with it. This experiment will investigate the impact of changing the dependency between the noise variables, $\epsilon(0)$ and $\epsilon(1)$, of $Y(0)$ and $Y(1)$, respectively, on the different conformal frameworks. The only difference with the previous experiment is that the noise correlation coefficient $c$ changes from $-1$ to $1$ in increments of $0.1$. The results are shown in Figure 11, 15, and 14.
@@
Setup A simulates a difficult nuisance component and an easy treatment effect. $d$ is the number of covariates and $n$ the number of samples. $X$ is created as an independent variable, $n$ rows and $d$ columns thus consisting of $[X_{0},X_{1},...,X_{d-1}]$ each with length $n$. The other variables are the expected outcome $b$, the propensity of receiving treatment $\pi$, $\tau$ the CATE, $W$ the treatment assignment, $Y(0)$ outcome for control, $Y(1)$ for treatment, $\epsilon(0)$ noise variable for control, $\epsilon(1)$ noise variable for treatment, and $Y$ the true outcome, all with length $n$.
@@
$\epsilon(0)\sim\mathcal{N}(0,\sigma^{2})$ (42)
$\epsilon(1)=c\epsilon(0)+(1-c)\mathcal{N}(0,\sigma^{2})$ (43)
$Y(0)=b-0.5\tau+\epsilon(0)$ (44)
$Y(1)=b+0.5\tau+\epsilon(1)$ (45)
$Y=Y(0)(1-W)+Y(1)W$ (46)
@@ G.3 Semi-synthetic experiments
The learners are also evaluated on semi-synthetic datasets. Four semi-synthetic datasets/experiments are used: IHDP [Hill, 2011], NLSM [Carvalho et al., 2019], EDU [Zhou et al., 2021], and the ACIC 2016[Dorie et al., 2019]. The simulations run 100 times with $\alpha=0.1$, 100 MC samples, and with split WCPS using the residual as a conformity measure. Since these datasets do not provide propensity scores, we perform logistic regression to estimate them. For the ACIC 2016 experiments, we run the simulations 10 times because it has 77 different settings. The results are found in Figure 2, 8, 9, 12, and 17.

### v6

@@ Appendix D The Epsilon Problem: (un)identifiability of ITE predictive distribution
We formally state the identifiability constraint: {theobox}
@@
Case 3. No identified copula If neither the independence nor a known copula is assumed, then infinitely many different copulas are compatible with the same pair of marginals but induce different distributions of the difference Y(1)Y(0)Y(1)-Y(0).
@@ D.1 Prediction intervals and violations of the noise assumption
$\mathbf{\boldsymbol{\varepsilon(0)\not\!\perp\!\!\!\perp\varepsilon(1)}}$: When the two noise distributions exhibit any dependency, inferring ITEs is challenging as we lack observations on the relationship between the distributions. If the noise distributions are positively correlated, potentially due to shared components in the outcome function and measurement errors in covariates, CCT- and CMC meta-learners could never guarantee valid predictive distributions. However, these models still retain the capability to yield conservatively valid prediction intervals.
@@ E.1 Probabilistic calibration of weighted conformal predictive system
The simplification at ()(*) can be made because we consider only a special case of weighted exchangeability, the covariate shift, where $w_{1:n}\equiv 1$ and $w_{n+1}((x,y))=w(x)=\frac{d\tilde{P}_{X}(x)}{dP(x)}$. Nevertheless, our findings can be readily applied to the general concept of weighted exchangeability.
@@ E.3 Existence universal consistent CCT- and CMC T-learner
$\int fdQ_{n}-\mathbb{E}_{\mathcal{P}}(f|X_{n+1})\rightarrow 0\qquad(n\rightarrow\infty)$ (22)
@@
the outcome variables are independent conditionally on the covariate set $X$. Then, by the definition of convolution and the fact that ITE=Y(1)Y(0)ITE=Y(1)-Y(0), the convolution of the true conditional cumulative distribution of outcome under treatment $F_{Y(1)|X}$ and outcome under negative control $F_{-Y(0)|X}$ equals the true conditional cumulative ITE distribution,
$F_{ITE|X}(y)=\int F_{Y(1)|X}(y+t)dF_{Y(0)|X}(t).$ (25)
@@ G.2 Synthetic experiments
This data generation process entails both Setup A and Setup B. For Setup A, where $\gamma=1$, represents an experiment where the treatment has no effect, while for Setup B, where $\gamma=0$, represents a heterogeneous treatment effect. X is sampled from a uniform distribution, $d$ is the number of covariates, $\tau$ represents the treatment effect, $\gamma$ is the treatment heterogeneity, $\pi$ is the propensity, $\varepsilon(0)$ and $\varepsilon(1)$ are Gaussian noise terms, and $W$ is the treatment assignment.
@@
$\varepsilon(1)\sim\mathcal{N}(0,1)$ (30)
$\varepsilon(0)\sim\mathcal{N}(0,1)$ (31)
$Y(0)=\tau_{0}+\varepsilon(0)$ (32)
$Y(1)=\tau+\varepsilon(1)$ (33)
@@ G.2.2 Nie and Wager synthetic data experiments
Besides the aforementioned experiments with the data-generating processes proposed by Nie and Wager [2021], we also perform a noise dependency impact experiment with it. This experiment will investigate the impact of changing the dependency between the noise variables, $\varepsilon(0)$ and $\varepsilon(1)$, of $Y(0)$ and $Y(1)$, respectively, on the different conformal frameworks. The only difference with the previous experiment is that the noise correlation coefficient $c$ changes from 1-1 to $1$ in increments of $0.1$. The results are shown in Figure 11, 15, and 14.
@@
Setup A simulates a difficult nuisance component and an easy treatment effect. $d$ is the number of covariates and $n$ the number of samples. $X$ is created as an independent variable, $n$ rows and $d$ columns thus consisting of $X_{0},X_{1},...,X_{d-1}]$ each with length $n$. The other variables are the expected outcome $b$, the propensity of receiving treatment $\pi$, $\tau$ the CATE, $W$ the treatment assignment, $Y(0)$ outcome for control, $Y(1)$ for treatment, $\varepsilon(0)$ noise variable for control, $\varepsilon(1)$ noise variable for treatment, and $Y$ the true outcome, all with length $n$.
@@
$\varepsilon(0)\sim\mathcal{N}(0,\sigma^{2})$ (42)
$\varepsilon(1)=c\varepsilon(0)+(1-c)\mathcal{N}(0,\sigma^{2})$ (43)
$Y(0)=b-0.5\tau+\varepsilon(0)$ (44)
$Y(1)=b+0.5\tau+\varepsilon(1)$ (45)
Y=Y(0)(1W)+Y(1)WY=Y(0)(1-W)+Y(1)W (46)
@@ G.3 Semi-synthetic experiments
The learners are also evaluated on semi-synthetic datasets. Four semi-synthetic datasets/experiments are used: IHDP [Hill, 2011], NLSM [Carvalho et al., 2019], EDU [Zhou et al., 2021], and the ACIC 2016[Dorie et al., 2019]. The simulations run 100 times with $\alpha=0.1$, 100 MC samples, and with split WCPS using the residual as a conformity measure. Since these datasets do not provide propensity scores, we perform logistic regression to estimate them. For the ACIC 2016 experiments, we run the simulations 10 times because it has 77 different settings. The results are found in Figure LABEL:fig:semi-synthetic-dist, 8, 9, 12, and 17.